Fermi National Laboratory


Accelerator Update for November 11, 2005 - November 14, 2005

Friday November 11

The day shift began with Operations monitoring store 4495 and the Recycler stash, with the Antiproton Source (Pbar) stacking, with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beam, with MiniBooNE, NuMI, and the Meson MCenter experiments taking beam.

At 8:17 AM, Operations reported a smoke alarm for the Tevatron (TeV) sector B0 compressor building. A cryo system technician said that a compressor had failed and discharged smoke and oil. The Fermi Fire Department, operators, and cryo experts responded.

Operations began preparing the Recycler to stash antiprotons at 9:42 AM.

Operations halted Linac beam for some necessary maintenance work at 10:20 AM.

The stash ended at 10:33 AM.

Interlock techs began testing the Booster safety system around 10:34 AM.

A Preaccelerator expert accessed the I- Source at 10:40 AM to replace its generator brushes. This job took forty minutes.

The Booster safety system test ended at 11:28 AM.

Pbar resumed stacking at 11:51 AM. Operations also resumed sending beam to MiniBooNE, NuMI, MTest, and MCenter.

Pbar experts began a reverse proton study at 12:49 PM.

An expert took the AP0 vault key while techs worked on the area’s air conditioning. This halted stacking at 12:53 PM. The worked ended and he returned the key about two hours later.

A Main Injector (MI) expert began a study period around 1:10 PM and ended at 6:22 PM.

Pbar reverse proton studies ended and Pbar resumed stacking at 7:14 PM.

The Recycler began preparing to stash at 10 PM.

At 10:12 PM, a Linac Klystron RF station (KRF6) tripped off and wouldn’t reset. Operations called in a Linac tech. This shut off all downstream beam.

The stash ended at 10:43 PM.

Saturday November 12

The midnight shift began with Operations monitoring store 4495 and the stash, with all other beam down due to KRF6 work.

At 12:59 AM, Linac techs took down KRF5 to compare with KRF6. Apparently, KRF6 was wired differently from the other stations.

The Linac techs completed their work on KRF6 at 3:30 AM.

Pbar resumed stacking at 3:34 AM.

Operations resumed sending beam to MCenter at 4:04 AM.

The Recycler began preparing to stash at 7:39 AM.

Operations terminated store 4495 at 8:05 AM.

The stash ended. Pbar, and MCenter resumed normal operations at 8:31 AM.

Operations prepared the D0 collision hall for an access at 8:37 AM. The experimenters came out at 10 AM.

Operations put the TeV through a dry squeeze at 10:38 AM.

Operations put the TeV into shot setup at 11 AM.

Operations established store 4497 at 12:54 PM with an initial luminosity of 148.95E30.

Recycler experts began a study period at 1:06 PM.

MTest and MCenter resumed taking beam at 1:33 PM.

The Recycler began preparing to stash at 3:10 PM. The stash ended. Pbar, MTest, and MCenter resumed normal operations at 4:09 PM.

The Recycler began preparing to stash at 7:37 PM. The stash ended and Pbar, MTest, and MCenter resumed normal operations by 8:26 PM.

At 11:09 PM, Operations reported that a Booster RF station (BRF12) tripped off for the first of four times in about two and half hours. It reset each time.

Sunday November 13

The midnight shift began with Operations monitoring store 4497 and the Recycler stash, with Pbar stacking, and with MiniBooNE, NuMI, MTest, and MCenter taking beam.

The Recycler began preparing to stash at 12:04 AM. The stash ended and Pbar, MTest, and MCenter resumed normal operations by 1:26 AM.

At 2:37 AM, the Linac debuncher tripped on a LCW fault. Operators called in an expert.

The Linac expert found a problem with the LCW pump configuration. Linac resumed sending beam at 4:13 AM.

The Recycler began preparing to stash at 6:35 AM. The stash ended and Pbar, MTest, and MCenter resumed normal operations by 7:20 AM.

At 11 AM, BRF12 tripped off on a ground fault twice in two-hours. It reset each time.

At 11:06 PM, MIRF8 and 18 tripped off for the first of two times in about three hours. It reset both times.

The Recycler began preparing to stash at 11:47 AM. The stash ended and Pbar, MTest, and MCenter resumed normal operations by 1:04 PM.

At 4:05 PM, BRF12 tripped off on a power fault indication for the first of five times in 3.5 hours. It reset each time.

Operations terminated store 4497 at 5:39 PM.

Operations put the TeV into shot setup at 5:53 PM.

At 6:40 PM, Operations reported that the TeV beam permit would not make up to TECAR. Operators investigated and replaced a bad heater firing unit.

Operations put the TeV through a dry squeeze at 7:46 PM.

Operations established store 4499 at 10:36 PM with an initial luminosity of 154.52E30.

The Recycler began preparing to stash at 10:47 PM. The stash ended and Pbar, MTest, and MCenter resumed normal operations by 11:47 PM.

Monday November 14

The midnight shift began with Operations monitoring store 4499 and the Recycler stash, with Pbar stacking, and with MiniBooNE, NuMI, MTest, and MCenter taking beam.

The Recycler began preparing to stash at 3:08 AM.

At 3:20 AM, a Recycler correction magnet power supply (CE504) tripped off. It reset okay. The Recycler didn’t lose any of its stash.

At 3:40 AM, a MiniBooNE power supply (E:H866) tripped off due to a ground fault indication. An operator called in an expert.

The stash ended. Pbar, MTest, and MCenter resumed normal operations by 4:18 AM.

The power supply experts reported at 6:35 AM that they had found and repaired a shorted cable on H866.

At 7:47 AM, Operations reported that H866 tripped off on a ground fault indication. Operations contacted the power supply experts. No beam to MiniBooNE.

The Plans for Monday

The plans for today are to stack, stash, and store. Interlock experts will conduct Pbar safety system checks today. Main Injector safety system checks will occur tomorrow.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 11/14/2005   email Fermilab